OpenStudy (kobeni-chan):

Can someone check my work please? Will medal It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (rainbow_dashie):

What do you think it is?

OpenStudy (kobeni-chan):

I set it up as 1/(x + 15) + 1/x = 1/4 Then I made the common denominator 4x(x + 15), and got 4 + 4(x+15) = x(x+15)

OpenStudy (kobeni-chan):

The 4 + 4(x + 15) = x(x + 15) was the numerators

OpenStudy (michele_laino):

sorry, if I call t_J time necessary to John to build the model car, then I got this equation fot t_J: \[t _{J}^{2}+7t _{J}-60=0\] from which \[t _{J}=5 hours,\] and t_B=time necessary to Brian to build the model car \[t _{B}=t _{J}+15=5+15=20hours\] is it right?

OpenStudy (michele_laino):

@kobeni-chan

OpenStudy (kobeni-chan):

oh I didn't think of it like that. In my class they were telling us to set them up as fractions, hold on a minute

OpenStudy (michele_laino):

@kobeni-chan your equation is right!

OpenStudy (michele_laino):

@kobeni-chan please, note that, it is the same that I used to find the above values!

OpenStudy (kobeni-chan):

ohh ok

OpenStudy (michele_laino):

in which your x, is my t_J

OpenStudy (kobeni-chan):

oh ok. Is It possible for you to figure out what x is through my way w/ the numerators? (sorry if that sounds bossy but I need to explain each step and that's the farthest I got)

OpenStudy (michele_laino):

ok! I will give you the entire explanation! @kobeni-chan

OpenStudy (anonymous):

Solve the following for b and j:\[b-15=j,\frac{1}{b}+\frac{1}{j}=\frac{1}{4} \]\[b=20,j=5 \]

OpenStudy (michele_laino):

I call W the wor and Brian have to do in order to build the car. Furthrtmore I call x the time that Jonn used to build the car, and y the time that Brian used to build the same car. Now from text of your problem we can write: y=x+15, do you agree? @kobeni-chan

OpenStudy (kobeni-chan):

ok yes

OpenStudy (michele_laino):

sorry ....I call with W the work that John and Brian have to do....

OpenStudy (michele_laino):

now rate r_J of John is: \[r _{J}=\frac{ W }{ x }\] and rate r_B of Brian is: \[r _{B}=\frac{ W }{ y }\] furthermore, from text of your problem, we can write: \[\frac{ W }{ x+y }=4\] do you agree?

OpenStudy (kobeni-chan):

Ok that makes sense, but my teacher wants me to use only x :/

OpenStudy (michele_laino):

ok! please note that: \[r _{B}=\frac{ W }{ x+15 }\] do you agree?

OpenStudy (kobeni-chan):

Yes

OpenStudy (michele_laino):

now, we can write, from text of your problem: \[r _{B}+r _{J}=\frac{ W }{ x+15 }+\frac{ W }{ x }=\frac{ W }{ 4 }\] do you agree?

OpenStudy (michele_laino):

@kobeni-chan

OpenStudy (kobeni-chan):

Ok yes

OpenStudy (michele_laino):

perfect! please divide by W, both sides of the above equation, and you will get your equation

OpenStudy (kobeni-chan):

I thought there wasn't a value for W? :/ sorry I feel confused

OpenStudy (kobeni-chan):

@Michele_Laino thanks for taking the time to explain all that to me :)

OpenStudy (michele_laino):

@kobeni-chan thank you!